# DGNS: Deformable Gaussian Splatting and Dynamic Neural Surface for Monocular Dynamic 3D Reconstruction

### v2

@@ 4.1 DNS for dynamic surface reconstruction
4.2 DGS for dynamic view synthesis
In contrast to existing methods using the 3DGS to model dynamic scenes [61, 29, 60, 24], we add the surface-aware density control, concentrating deformable 3D Gaussian points around the surface area, which enhances the model’s ability to capture both the geometry and color of dynamic surfaces. Additionally, we incorporate surface normal supervision using normals derived from foundation models. These components are detailed below.
Surface-aware density control: Utilizing the object’s surface can serve as effective guidance for positioning Gaussian primitives to enhance rendering quality [52, 64, 32]. However, directly aligning Gaussian primitives to the surface often causes a decline in rendering quality [66, 18]. To address this, we adopt a surface-aware density control strategy, similar to GSDF [64] for static scenes, to optimize the distribution of Gaussian primitives. Specifically, the zero-level set of DNS (Eq.3) in observation space is used to guide Gaussian growth (split/clone) and pruning operations. Gradient-based adaptive density control[23] and the SDF values of Gaussian primitives from the DNS module are employed to fine-tune Gaussian placement and density. For each Gaussian primitive $\mathbf{x}_{g}$ in DGS canonical space, we determine its location in observation space as ($\mathbf{x}_{g}+\delta\mathbf{x}_{g}$), where $(\delta\mathbf{x}_{g},\delta\mathbf{r}_{g},\delta\mathbf{s}_{g})=F_{\theta}(\gamma(\mathbf{x}_{g}),\gamma(t))$ through the GS deformation field. The SDF distance of $\mathbf{x}_{g}$ is then calculated as ${d}_{g}=\mathcal{F}(\mathcal{H}(\mathbf{x}{g}+\delta\mathbf{x}_{g},t))$ Accordingly, the criteria for Gaussian growth are defined as follows:
$\epsilon_{g}=\nabla_{\mathbf{x}_{g}}+w_{g}\phi(d_{g}),$ (8)
@@ 4.2 DGS for dynamic view synthesis
$\epsilon_{p}=\sigma_{p}-w_{p}(1-\phi(d_{g}),$ (9)
where $\sigma_{p}$ is the sum of opacity across $K$ iterations and the weighting parameter $w_{p}$ controls the influence of the SDF distance on running criteria. Gaussian primitives will be removed if $\epsilon^{p}$ is below a set threshold $\tau^{p}$.
Normal supervision: To ease the under-constrained optimization problem of monocular dynamic scene reconstruction, we add the normal loss to regularize the GS module. The normal direction of each Gaussian primitive can be approximated with the direction of the axis with the minimum scaling factor [8]. The normal in the world coordinate system is defined as: $n=\mathbf{R}[k,:]\in\mathbb{R}^{3},\hskip 2.5ptk=\arg\min([s_{1},s_{2},s_{3}])$, where $s_{1},s_{2},s_{3}$ are the Gaussian scales and $\mathbf{R}$ is the Gaussian rotation matrix. Similar to rendering color Eq. 1, the normal vector of a point $p$ in the screen space can be rendered as $\hat{N}^{g}(p)=\sum_{i\in N}n_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$ Therefore, the normal regularization loss can be calculated as:
@@
4.3 Optimization
For the GS module, the image loss measures the difference between the rendered RGB images and ground truth images and usually includes two rendering losses $\mathcal{L}_{1}$ and $\mathcal{L}_{\text{ssim}}$, supplemented by a normal loss.
@@ 4.3 Optimization
where $\lambda_{I}$ and $\lambda_{\text{gn}}$ are the weighting coefficients. Similarly, the image loss in the DNS module is supervised by the $\mathcal{L}_{1}$ loss. The SDF will be regularized by Eikonal loss $\mathcal{L}_{eik}$ [17]. Moreover, the SDF is supervised by the filtered point which generated the $\mathcal{L}_{\text{sdf}}$ as Equ 6.
@@
where $\lambda_{\text{sdf}}$, $\lambda_{\text{nn}}$, and $\lambda_{\text{eik}}$ are the weighting parameters for each loss term. The final total loss will be
@@
5 Experiments
5.1 Setup
Method Duck Horse Bird CD $\downarrow$ EMD $\downarrow$ PSNR $\uparrow$ CD $\downarrow$ EMD $\downarrow$ PSNR $\uparrow$ CD $\downarrow$ EMD $\downarrow$ PSNR $\uparrow$ D-NeRF 0.934 0.073 29.79 1.685 0.280 25.47 1.532 0.163 23.85 NDR 2.235 0.083 24.79 0.359 0.148 27.23 0.444 0.130 25.43 K-Plane 1.085 0.055 33.36 1.480 0.239 28.11 0.742 0.131 23.72 HexPlane 2.161 0.090 32.11 1.750 0.199 26.78 4.158 0.178 22.19 TiNeuVox 0.969 0.059 34.33 1.918 0.246 28.16 8.264 0.215 25.55 D-3DGS 1.643 0.1146 36.76 0.941 0.203 35.66 1.398 0.138 29.38 Dg-mesh 0.790 0.047 32.89 0.299 0.168 27.10 0.557 0.128 22.98 GNS2(ours) 0.773 0.046 37.39 0.289 0.142 35.52 0.440 0.126 30.23 Method Beagle Torus2sphere Girlwalk CD $\downarrow$ EMD $\downarrow$ PSNR $\uparrow$ CD $\downarrow$ EMD $\downarrow$ PSNR $\uparrow$ CD $\downarrow$ EMD $\downarrow$ PSNR $\uparrow$ D-NeRF 1.001 0.149 34.47 1.760 0.250 24.23 0.601 0.190 28.63 NDR 0.747 0.106 28.39 1.767 0.176 25.19 0.482 0.158 29.98 K-Plane 0.810 0.122 38.33 1.793 0.161 31.22 0.495 0.173 32.12 HexPlane 0.870 0.115 38.03 2.190 0.190 29.71 0.597 0.155 31.77 TiNeuVox 0.874 0.129 38.97 2.115 0.203 28.76 0.568 0.184 32.81 D-3DGS 0.898 0.211 38.56 1.487 0.193 30.22 0.883 0.270 38.75 Dg-mesh 0.639 0.117 34.53 1.607 0.172 26.61 0.726 0.136 28.64 GNS2(ours) 0.643 0.101 39.29 0.456 0.153 29.76 0.413 0.124 38.64
@@ 5.1 Setup
Table 1: Mesh reconstruction and rendering quality results of our method compared with other baselines. Reconstructed meshes are measured with Chamfer Distance (CD) and Earth Mover Distance (EMD) with the ground truth mesh. Rendering quality is measured with Peak Signal-to-Noise Ratio (PSNR). The color of each cell indicates the best, second, and thirds score, and the third best results. In general, our method produces a better reconstruction and rendering quality among all metrics.
@@
Dataset and baseline: In the main body of this work, we conducted experimental evaluations on two public monocular video datasets (i.e. D-NeRF [43] and Dg-mesh [29]). Evaluations on other monocular video datasets, such as Nerfies [41], are included in the Supplementary material. D-NeRF includes eight sets of dynamic scenes featuring complex motion, such as articulated objects and human actions. Dg-mesh provides six sets of dynamic scenes with ground truth for each object’s deformable 3D structure. Both datasets have images at an 800$\times$800 resolution, with 100 to 200 images per scene. To demonstrate the effectiveness of our method, we compared it with seven baselines: D-NeRF [43], NDR [3], K-Plane [13], HexPlane [5], TiNeuVox [10], Dg-mesh [29], and D3DGS [61]. D-NeRF, TiNeuVox, and NDR use implicit representations with a deformation field to map the dynamic observation space to static canonical space, while K-Plane and HexPlane employ 4D feature volumes with volume factorization. In contrast, Dg-mesh and D3DGS rely on explicit 3D Gaussian representations to model dynamic scenes. Our method integrates SDF and 3D Gaussians, achieving state-of-the-art performance in both 3D reconstruction and view synthesis.
[…]
Figure 4: Qualitative comparison on the Dg-mesh dataset. The samples, from top to bottom, are Horse, Torus2sphere, Bird, and Beagle. Compared to other baselines, our results are the closest to the ground truth (GT).
5.2 Results
Dg-mesh dataset: The Dg-mesh dataset [29] provides geometrical ground truth for dynamic objects at each timeframe, enabling a thorough quantitative evaluation of our method, DGNS, for both 3D reconstruction and novel-view synthesis tasks across six objects. As shown in Tab 1, our method consistently achieves the lowest CD and EMD scores across nearly all object categories. This performance highlights DGNS’s ability to capture fine-grained dynamic spatial structures with higher accuracy than competing methods, while also surpassing baselines in rendering quality. In terms of reconstruction accuracy, Dg-mesh frequently ranks second, especially in achieving lower CD and EMD scores, yet this comes with a noticeable compromise in rendering quality. Conversely, while D3DGS exhibits competitive performance in novel-view synthesis, it fails to achieve the same low CD and EMD scores necessary for high-accuracy 3D reconstruction. Our method is unique in offering consistently superior performance across both 3D reconstruction and rendering quality, providing a comprehensive solution for tasks requiring both structural precision and visual fidelity.
The qualitative results, presented in Fig 4, further substantiate these findings, showing that meshes reconstructed by DGNS most closely resemble the ground truth. For example, Gaussian-based methods like D3DGS [61] struggle with surface accuracy due to floating Gaussian points, resulting in less cohesive 3D surfaces. While Dg-mesh [29] improves upon this with an anchoring process to reduce floating points, the resultant surface still lacks sufficient detail and smoothness. The NDR [3] method encounters challenges with specific object parts, such as accurately reconstructing the legs in Girlwalk, the tails of Bird, and the inner surface of Torus2sphere.
Figure 5: Qualitative comparison on the D-NeRF dataset. The samples from right to left are Hellwarrior and Standup. Our method can achieve 3D reconstructions with smooth surfaces.
D-NeRF dataset: The comparison of our method against baselines is presented in Tab 2, demonstrating that our method, DGNS, provides a robust balance across PSNR, SSIM, and LPIPS metrics across various scenes. This balance underscores its effectiveness in novel-view synthesis tasks. While D3DG slightly outperforms our method in rendering quality across some scenes, DGNS consistently surpasses other baselines, including Dg-mesh, the method previously noted for its efficacy in 3D reconstruction.
Notably, in the D-NeRF dataset, mesh ground truths are unavailable. Thus, we have supplemented the quantitative analysis with a qualitative comparison, shown in Fig 5. The qualitative assessment reveals that our method yields superior performance in 3D reconstruction fidelity, even when visually assessed against high-performing baselines. Although D3DGS achieves top scores in rendering quality, it does not maintain the structural accuracy in 3D reconstruction that our method consistently delivers. In contrast, our method demonstrates state-of-the-art rendering quality while excelling in structural reconstruction, presenting a compelling solution for applications requiring both high fidelity in visual output and accurate 3D geometry. More qualitative results are included in the Supplementary material.
Method Hell Warrior Mutant Hook Bouncing Balls PSNR ↑ SSIM ↑ LPIPS ↓ PSNR ↑ SSIM ↑ LPIPS ↓ PSNR ↑ SSIM ↑ LPIPS ↓ PSNR ↑ SSIM ↑ LPIPS ↓ D-NeRF 24.06 0.944 0.071 30.31 0.967 0.039 29.02 0.960 0.055 38.17 0.989 0.032 NDR 31.58 0.971 0.035 29.83 0.958 0.049 26.62 0.952 0.071 36.34 0.986 0.039 TiNeuVox 27.10 0.964 0.077 31.87 0.961 0.047 30.61 0.960 0.059 40.23 0.993 0.044 Tensor4D 31.26 0.975 0.024 29.11 0.965 0.056 24.47 0.943 0.064 26.56 0.992 0.044 K-Planes 24.58 0.952 0.084 32.50 0.951 0.066 28.12 0.949 0.052 40.05 0.993 0.032 D-3DG 41.54 0.987 0.023 42.63 0.995 0.005 37.42 0.987 0.014 41.01 0.995 0.009 Dg-mesh 25.46 0.959 0.074 30.40 0.968 0.055 27.88 0.954 0.074 29.15 0.969 0.099 GNS2(ours) 41.68 0.988 0.023 41.47 0.984 0.011 36.34 0.979 0.023 40.71 0.994 0.012 Method Lego T-Rex Stand Up Jumping Jacks PSNR ↑ SSIM ↑ LPIPS ↓ PSNR ↑ SSIM ↑ LPIPS ↓ PSNR ↑ SSIM ↑ LPIPS ↓ PSNR ↑ SSIM ↑ LPIPS ↓ D-NeRF 25.56 0.936 0.082 30.61 0.967 0.054 33.13 0.978 0.036 32.70 0.978 0.039 NDR 24.54 0.936 0.084 23.47 0.963 0.047 30.22 0.961 0.047 29.02 0.957 0.043 TiNeuVox 26.64 0.953 0.078 31.25 0.967 0.048 34.61 0.980 0.033 33.49 0.977 0.041 Tensor4D 23.24 0.942 0.089 23.86 0.954 0.054 26.30 0.938 0.056 24.20 0.925 0.067 K-Planes 28.91 0.970 0.035 30.43 0.970 0.031 33.10 0.979 0.031 31.11 0.971 0.047 D-3DG 33.07 0.979 0.018 38.10 0.993 0.010 44.62 0.995 0.006 37.72 0.990 0.013 Dg-mesh 21.29 0.838 0.159 28.95 0.959 0.065 30.21 0.974 0.051 31.77 0.980 0.045 GNS2(ours) 29.48 0.974 0.026 37.68 0.989 0.012 44.81 0.996 0.007 38.60 0.993 0.011
@@ 5.2 Results
Table 2: Rendering quality results of our method compared with other baselines. Our method is closely aligned with the state-of-the-art D-3DGS, yet significantly outperforms other baseline methods.
5.3 Ablation study

### v3

@@ 4.1. DNS for Dynamic Surface Reconstruction
4.2. DGS for Dynamic View Synthesis
In contrast to existing methods using the 3DGS to model dynamic scenes (Yang et al., 2024; Liu et al., 2024; Yang et al., 2023a; Kratimenos et al., 2023), we add the surface-aware density control with the geometry guidance from DNS module (as shown with green arrow in fig. 2), concentrating deformable 3D Gaussian points around the surface area, which enhances the model’s ability to capture both the geometry and color of dynamic surfaces. Additionally, we incorporate surface normal supervision using normals derived from foundation models. These components are detailed below.
Surface-aware Density Control. Utilizing the object’s surface can serve as effective guidance for positioning Gaussian primitives to enhance rendering quality (Wang et al., 2023; Yu et al., 2024; Lu et al., 2024). However, directly aligning Gaussian primitives to the surface often causes a decline in rendering quality (Zhang et al., 2024a; Guédon and Lepetit, 2024). To address this, we adopt a surface-aware density control strategy, similar to GSDF (Yu et al., 2024) for static scenes, to optimize the distribution of Gaussian primitives. Specifically, the zero-level set of DNS (see Eq. (3)) in observation space is used to guide Gaussian growth (split/clone) and pruning operations. Gradient-based adaptive density control(Kerbl et al., 2023) and the SDF values of Gaussian primitives from the DNS module are employed to fine-tune Gaussian placement and density. For each Gaussian primitive $\mathbf{x}_{g}$ in DGS canonical space, we determine its location in observation space as ($\mathbf{x}_{g}+\delta\mathbf{x}_{g}$), where $(\delta\mathbf{x}_{g},\delta\mathbf{r}_{g},\delta\mathbf{s}_{g})=F_{\theta}(\gamma(\mathbf{x}_{g}),\gamma(t))$ through the GS deformation field. The SDF distance of $\mathbf{x}_{g}$ is then calculated as ${d}_{g}=\mathcal{F}(\mathcal{H}(\mathbf{x}{g}+\delta\mathbf{x}_{g},t))$ . Accordingly, the criteria for Gaussian growth are defined as follows:
(8) $\epsilon_{g}=\nabla_{\mathbf{x}_{g}}+w_{g}\phi(d_{g}),$
@@ 4.2. DGS for Dynamic View Synthesis
(9) $\epsilon_{p}=\sigma_{p}-w_{p}(1-\phi(d_{g})),$
where $\sigma_{p}$ is the sum of opacity across $K$ iterations and the weighting parameter $w_{p}$ controls the influence of the SDF value. Gaussian primitives will be removed if $\epsilon^{p}$ is below a set threshold $\tau^{p}$.
Normal Supervision. To ease the under-constrained optimization problem of monocular dynamic scene reconstruction, we add the normal loss to regularize the GS module. The normal direction of each Gaussian primitive can be approximated with the direction of the axis with the minimum scaling factor (Chen et al., 2023). The normal in the world coordinate system is defined as: $n=\mathbf{R}[k,:]\in\mathbb{R}^{3},\hskip 2.5ptk=\arg\min([s_{1},s_{2},s_{3}])$, where $s_{1},s_{2},s_{3}$ are the Gaussian scales and $\mathbf{R}$ is the Gaussian rotation matrix. Similar to rendering color in eq. 1, the normal vector of a point $p$ in the screen space can be rendered as $\hat{N}^{g}(p)=\sum_{i\in N}n_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$ . Therefore, the normal regularization loss can be calculated as:
@@
Figure 4. Qualitative comparison on the Dg-mesh dataset. The samples, from top to bottom, are Horse, Torus2sphere, Bird, and Beagle. Compared to other baselines, our results are the closest to the ground truth (GT).
4.3. Optimization
For the GS module, the image loss measures the difference between the rendered RGB images and ground truth images. Usually, it includes two rendering losses $\mathcal{L}_{1}$ and $\mathcal{L}_{\text{ssim}}$, supplemented by a normal loss:
@@ 4.3. Optimization
where $\lambda_{I}$ and $\lambda_{\text{gn}}$ are the weighting coefficients. Similarly, the image loss in the DNS module is supervised by the $\mathcal{L}_{1}$ loss. The SDF will be regularized by Eikonal loss $\mathcal{L}_{eik}$ (Gropp et al., 2020). Moreover, the SDF is supervised by the filtered point which generated the $\mathcal{L}_{\text{sdf}}$ as Eq. (6). Thus, the loss is designed as:
@@
where $\lambda_{\text{sdf}}$, $\lambda_{\text{nn}}$, and $\lambda_{\text{eik}}$ are the weighting parameters for each loss term. The final total loss will be:
@@
With the $\mathcal{L}$, the proposed DGNS learns hybrid representations jointly across two tasks (i.e., 3D geometry reconstruction and novel view synthesis) using a unified framework, and two modules mutually benefit from each other through efficient ray-sampling, depth regularization, and guided density control.
5. Experiments
5.1. Setup
@@
Table 1. Mesh reconstruction and rendering quality results of our method compared to other baselines on Dg-mesh (Liu et al., 2024). Reconstructed meshes are measured with Chamfer Distance (CD) and Earth Mover Distance (EMD) with the ground truth mesh. Rendering quality is measured with Peak Signal-to-Noise Ratio (PSNR). The color of each cell indicates the best, second, and third scores, and the third-best results. In general, our method produces a better reconstruction and rendering quality.
Method Duck Horse Bird CD $\downarrow$ EMD $\downarrow$ PSNR $\uparrow$ CD $\downarrow$ EMD $\downarrow$ PSNR $\uparrow$ CD $\downarrow$ EMD $\downarrow$ PSNR $\uparrow$ D-NeRF 0.934 0.073 29.79 1.685 0.280 25.47 1.532 0.163 23.85 NDR 2.235 0.083 24.79 0.359 0.148 27.23 0.444 0.130 25.43 K-Plane 1.085 0.055 33.36 1.480 0.239 28.11 0.742 0.131 23.72 HexPlane 2.161 0.090 32.11 1.750 0.199 26.78 4.158 0.178 22.19 TiNeuVox 0.969 0.059 34.33 1.918 0.246 28.16 8.264 0.215 25.55 D3DGS 1.643 0.1146 36.76 0.941 0.203 35.66 1.398 0.138 29.38 Dg-mesh 0.790 0.047 32.89 0.299 0.168 27.10 0.557 0.128 22.98 DGNS (ours) 0.773 0.046 37.39 0.289 0.142 35.52 0.440 0.126 30.23 Method Beagle Torus2sphere Girlwalk CD $\downarrow$ EMD $\downarrow$ PSNR $\uparrow$ CD $\downarrow$ EMD $\downarrow$ PSNR $\uparrow$ CD $\downarrow$ EMD $\downarrow$ PSNR $\uparrow$ D-NeRF 1.001 0.149 34.47 1.760 0.250 24.23 0.601 0.190 28.63 NDR 0.747 0.106 28.39 1.767 0.176 25.19 0.482 0.158 29.98 K-Plane 0.810 0.122 38.33 1.793 0.161 31.22 0.495 0.173 32.12 HexPlane 0.870 0.115 38.03 2.190 0.190 29.71 0.597 0.155 31.77 TiNeuVox 0.874 0.129 38.97 2.115 0.203 28.76 0.568 0.184 32.81 D3DGS 0.898 0.211 38.56 1.487 0.193 30.22 0.883 0.270 38.75 Dg-mesh 0.639 0.117 34.53 1.607 0.172 26.61 0.726 0.136 28.64 DGNS (ours) 0.643 0.101 39.29 0.456 0.153 29.76 0.413 0.124 38.64
Figure 5. Qualitative results demonstrating the temporal evolution of 3D meshes reconstructed by our method. Rows (from top to bottom) show the Torus2sphere, Horse, and Bird sequences from the dg-mesh dataset. Columns (from left to right) depict different time points
@@
Dataset and Baseline. In this work, we conducted experimental evaluations on three public monocular video datasets: two synthetic datasets, D-NeRF (Pumarola et al., 2021) and Dg-mesh (Liu et al., 2024), and one real dataset, Nerfies (Park et al., 2021). Comprehensive quantitative analyses and selected qualitative results are presented in the main paper, while additional qualitative results can be found in the supplementary material. D-NeRF includes eight sets of dynamic scenes featuring complex motion, such as articulated objects and human actions. Dg-mesh provides six sets of dynamic scenes with ground truth for each object’s deformable 3D structure. Both datasets have images at $800\times 800$ resolution, with $100$ to $200$ images per scene. To demonstrate the effectiveness of our method, we compared it with seven baselines: D-NeRF (Pumarola et al., 2021), NDR (Cai et al., 2022), K-Plane (Fridovich-Keil et al., 2023), HexPlane (Cao and Johnson, 2023), TiNeuVox (Fang et al., 2022), Dg-mesh (Liu et al., 2024), and D3DGS (Yang et al., 2024). D-NeRF, TiNeuVox, and NDR use implicit representations with a deformation field to map the dynamic observation space to static canonical space, while K-Plane and HexPlane employ 4D feature volumes with volume factorization. In contrast, Dg-mesh and D3DGS rely on explicit 3D Gaussian representations to model dynamic scenes. Our method integrates SDF and 3D Gaussians, achieving state-of-the-art performance in both 3D reconstruction and view synthesis.
Implementations.
For the DNS module, we utilize the bijective mapping network (Dinh et al., 2016; Cai et al., 2022) to learn the deformation field to mapping points from observation space back to canonical space, and canonical space is the hybrid of hash-grid encoders and 4-layer MLP for speedup. The entire training process includes a warm-up phase (0 to 1010k iterations) followed by joint training (1010k to 4040k iterations). For efficient ray sampling, the scaling factor $s$ is set to $3$ from 1010k to 2020k iterations for coarse search, and then to $1$ for the remainder of the training. The DGS model uses an 8-layer MLP (256 channels) for deformation learning. DGS takes around 1010k iterations to make an accurate depth prediction, from which depth guidance for ray and SDF supervision in NS starts. After the 1515k iteration warm-up, the geometry guidance for density control in DGS begins from 1515k iterations. Joint optimization lasts for 2525k iterations. The entire training process includes a warm-up phase (0 to 1515k iterations) followed by joint training (1515k to 4040k iterations). The weight $\lambda_{I}$ is set to $0.8$ for calculating image loss, while $\lambda_{gn}$ starts to take effect at 1010k iterations with a value of $0.1$. The deformable network activates after 33k iterations, with density control starting at 500 iterations. Gaustudio (Ye et al., 2024) extracts 3D meshes from dynamic Gaussian primitives (Yang et al., 2024). All experiments used NVIDIA RTX A6000 GPU (48GB).
5.2. Results
@@
Table 2. Rendering quality results of our method compared with other baselines on D-NeRF (Pumarola et al., 2021). Our method is on par with the state-of-the-art D3DGS, yet significantly outperforms other baseline methods.
Method Hell Warrior Mutant Hook Bouncing Balls PSNR ↑ SSIM ↑ LPIPS ↓ PSNR ↑ SSIM ↑ LPIPS ↓ PSNR ↑ SSIM ↑ LPIPS ↓ PSNR ↑ SSIM ↑ LPIPS ↓ D-NeRF 24.06 0.944 0.071 30.31 0.967 0.039 29.02 0.960 0.055 38.17 0.989 0.032 NDR 31.58 0.971 0.035 29.83 0.958 0.049 26.62 0.952 0.071 36.34 0.986 0.039 TiNeuVox 27.10 0.964 0.077 31.87 0.961 0.047 30.61 0.960 0.059 40.23 0.993 0.044 Tensor4D 31.26 0.975 0.024 29.11 0.965 0.056 24.47 0.943 0.064 26.56 0.992 0.044 K-Planes 24.58 0.952 0.084 32.50 0.951 0.066 28.12 0.949 0.052 40.05 0.993 0.032 D3DGS 41.54 0.987 0.023 42.63 0.995 0.005 37.42 0.987 0.014 41.01 0.995 0.009 Dg-mesh 25.46 0.959 0.074 30.40 0.968 0.055 27.88 0.954 0.074 29.15 0.969 0.099 DGNS (ours) 41.68 0.988 0.023 41.47 0.984 0.011 36.34 0.979 0.023 40.71 0.994 0.012 Method Lego T-Rex Stand Up Jumping Jacks PSNR ↑ SSIM ↑ LPIPS ↓ PSNR ↑ SSIM ↑ LPIPS ↓ PSNR ↑ SSIM ↑ LPIPS ↓ PSNR ↑ SSIM ↑ LPIPS ↓ D-NeRF 25.56 0.936 0.082 30.61 0.967 0.054 33.13 0.978 0.036 32.70 0.978 0.039 NDR 24.54 0.936 0.084 23.47 0.963 0.047 30.22 0.961 0.047 29.02 0.957 0.043 TiNeuVox 26.64 0.953 0.078 31.25 0.967 0.048 34.61 0.980 0.033 33.49 0.977 0.041 Tensor4D 23.24 0.942 0.089 23.86 0.954 0.054 26.30 0.938 0.056 24.20 0.925 0.067 K-Planes 28.91 0.970 0.035 30.43 0.970 0.031 33.10 0.979 0.031 31.11 0.971 0.047 D3DGS 33.07 0.979 0.018 38.10 0.993 0.010 44.62 0.995 0.006 37.72 0.990 0.013 Dg-mesh 21.29 0.838 0.159 28.95 0.959 0.065 30.21 0.974 0.051 31.77 0.980 0.045 DGNS (ours) 29.48 0.974 0.026 37.68 0.989 0.012 44.81 0.996 0.007 38.60 0.993 0.011
Dg-mesh Dataset. The Dg-mesh dataset (Liu et al., 2024) provides geometrical ground truth for dynamic objects at each timeframe, enabling a thorough quantitative evaluation of our method, DGNS, for both 3D reconstruction and novel-view synthesis tasks across six objects. As shown in Tab. 1, our method consistently achieves the lowest CD and EMD scores across nearly all object categories. This performance highlights DGNS’s ability to capture fine-grained dynamic spatial structures with higher accuracy than competing methods, while also surpassing baselines in rendering quality. Regarding reconstruction accuracy, Dg-mesh frequently ranks second, especially in achieving lower CD and EMD scores, however, this improvement comes at the cost of rendering quality. Conversely, while D3DGS exhibits competitive performance in novel-view synthesis, it fails to achieve the same low CD and EMD scores necessary for high-accuracy 3D reconstruction. Our method is unique in offering consistently superior performance across both 3D reconstruction and rendering quality, providing a comprehensive solution for tasks requiring both structural precision and visual fidelity.
The qualitative results are presented in Figs. 4 and 5. In Fig. 4, it is shown that meshes reconstructed by DGNS (ours) most closely resemble the ground truth. For example, Gaussian-based methods like D3DGS (Yang et al., 2024) struggle with surface accuracy due to floating Gaussian points, resulting in less cohesive 3D surfaces. While Dg-mesh (Liu et al., 2024) improves upon this with an anchoring process to reduce floating points, the resultant surface still lacks sufficient detail and smoothness. The NDR (Cai et al., 2022) method encounters challenges with specific object parts, such as accurately reconstructing the legs in Girlwalk, the tails of Bird, and the inner surface of Torus2sphere. Fig. 5 demonstrates our method’s capability to effectively model dynamic samples exhibiting extreme deformations, rapid motion, or topological changes, which are critical for validating its robustness in real-world scenarios. Additional qualitative results are provided in the supplementary materials.
Figure 6. Qualitative comparison on the D-NeRF dataset. The samples from right to left are Hellwarrior and Standup. Our method can achieve 3D reconstructions with smooth surfaces.
D-NeRF Dataset. The comparison of our method against baselines is presented in Tab. 2, demonstrating that our method, DGNS, provides a robust balance across PSNR, SSIM, and LPIPS metrics across various scenes. This balance underscores its effectiveness in novel-view synthesis tasks. While our method achieves comparable performance with D3DGS in rendering quality across some scenes, DGNS consistently surpasses other baselines, including Dg-mesh, the method previously noted for its efficacy in 3D reconstruction.
Notably, mesh ground truths are unavailable in the D-NeRF dataset. Thus, we have supplemented the quantitative analysis with a qualitative comparison, shown in Fig. 6. The qualitative assessment reveals that our method yields superior performance in 3D reconstruction fidelity, even when visually assessed against high-performing baselines. Although D3DGS achieves top scores in rendering quality, it does not maintain the structural accuracy in 3D reconstruction that our method consistently delivers. In contrast, our method demonstrates state-of-the-art rendering quality while excelling in structural reconstruction, presenting a compelling solution for applications requiring high fidelity in visual output and accurate 3D geometry. More qualitative results are included in the Supplementary material.
Nerfies Dataset. Due to the absence of ground-truth data for the Nerfies dataset (Park et al., 2021), we provide a qualitative comparison, shown in Fig. 7. As illustrated, our method produces smoother surfaces compared to DG-mesh and achieves more accurate geometric reconstructions than NDR.
Figure 7. Qualitative comparison on the Nerfies dataset.
5.3. Ablation Study